OpenStudy (anonymous):

vector "a" makes equal angles withx,y,z axis.find the value of its components(in terms of magnitude of "a".

OpenStudy (anonymous):

If there is a vector translating the same angle with the three axis, then it will also have the same value proyected over each axis. Such vector will be sticking at the very center of its quadrant, same as vector "a"= Xi+Yj+Zk where in this case X=Y=Z so then let X,Y,Z be called W, the vector's magnitude is \[\left| a \right|=\sqrt{W^{2}+W^{2}+W^{2}}\]then\[\left| a \right|=\sqrt{3W^{2}}=W \sqrt{3}\]